# Adaptive Deep Learning for Multiclass Breast Cancer Classification via Misprediction Risk Analysis

### v1

@@ 4.2 Attention-based Risk Model Construction
The existing binary model supposes risk features are mutually independent. However, in real scenarios, risk features are to a large extent correlated, which is especially true in the multi-class scenario. Therefore, we present a new dynamic risk model with attention-based weighting for multi-class risk analysis. The attention mechanism dynamically adjusts the importance of risk features based on the context of each class, improving multiclass classification by weighting features according to their relevance [42, 26].
@@
In a multi-class classification task, consider an image $d_{i}$ that can belong to one of four classes: A, B, C, or D. In the learnable Gaussian model, each class has a set of mean ($\mu_{A},\mu_{B},\mu_{C},\mu_{D}$) and variance ($\sigma_{A}^{2},\sigma_{B}^{2},\sigma_{C}^{2},\sigma_{D}^{2}$) parameters that are learned from the risk features. These parameters are used to compute the Gaussian weights, however, even with these learnable parameters, the model still uses fixed, class-specific parameters that do not change dynamically based on the image content. In contrast, the attention mechanism assigns dynamic weights based on the risk features. For example, for image $d_{i}$, the attention mechanism might assign $\omega_{A}=0.6$, $\omega_{B}=0.1$, $\omega_{C}=0.2$, and $\omega_{D}=0.1$, emphasizing class A more than the others based on the image’s specific features. This dynamic adjustment allows the attention mechanism to focus on the most relevant classes for each input image, whereas the Gaussian model is constrained by pre-learned, static parameters.
@@ 4.3 Risk Model Training
Class imbalance is common in most histopathology image datasets, as shown in Tables 4 and 5, particularly in the training data. To address this issue, we propose to neutralize class bias in expectation and variance for balanced learning and prediction. The neutralization operation is formally defined by Equation 15, and 16 as follows:
@@
7:Maximum number of training iterations $N$
@@
where $\gamma$ denotes the risk score of the particular classes of corresponding images . Considering this scenario, the posterior probability $p_{ij,c}$ of the image $(d_{i},c_{i})$ is ranked higher than $(d_{j},c_{j})$ is given by:
@@
In a breast cancer classification task with four classes (benign, malignant, in situ, and invasive), consider two images, $d_{i}$ and $d_{j}$, with predicted class probability vectors $d_{i}=[0.2,0.6,0.1,0.1]$ and $d_{j}=[0.3,0.4,0.2,0.1]$. Performing pairwise comparisons, $d_{i}$ wins against $d_{j}$ in the malignant class (1 win) but loses in the benign and in situ classes (0 wins each), while both images tie in the invasive class. The total win count results in $W_{i}=1$ for $d_{i}$ and $W_{j}=2$ for $d_{j}$. Consequently, $d_{j}$ is ranked higher due to its superior win count, demonstrating the effectiveness of the voting-based mechanism for class-based image ranking.
@@
We also introduced dropout during training to randomly set a percentage of input units to zero. This prevents co-adaptation and promotes the learning of more robust features for each class. Dropout is inactive during evaluation, with activations scaled for consistency. In multiclass classification, dropout helps prevent overfitting by ensuring that the model does not overly rely on specific features. We set the dropout rate to 0.5 and fine-tuned it for optimal performance.
@@ 5 Solution for MultiRisk Based Adaptive Training
Neural networks using softmax for multiclass tasks often produce overly confident predictions, which can mislead training and impair generalization [13]. Class imbalance can further degrade performance on minority classes [20], and risk estimation may fail to accurately reflect misclassification likelihood, particularly in complex multiclass scenarios [19].
@@
  • 1.
    Traditional Pre-Training Phase: During this initial phase, a deep model undergoes training using conventional methods, relying on labeled training data. In our method, we use four CNN deep models and train all of them on the given workload with arguably the best parameters. The aim is to try to achieve the best test efficacy for each model.
  • 2.
    Risk Based Adaptive Training Phase: This phase aims to enhance model performance by focusing on high-risk instances and managing prediction uncertainties. This framework integrates temperature scaling to reduce overconfidence and Value at Risk (VaR) to assess misclassification risks. The goal is to improve the robustness and accuracy of the classifier. The step-wise process is given below:
Algorithm 3 Risk-Based Adaptive Training for Multiclass Breast Cancer Detection
Input: A workload of Multiclass Breast Cancer histopathological images $D$ consisting of $D_{s}$, $D_{v}$, $D_{t}$ and initial classifier $g(\omega)$;
@@
1:$\omega_{0}\leftarrow$ Initialize $\omega$ with random values;
2:for $k=0$ to $m-1$ do
@@
In the second phase, the algorithm first uses the validation set $D_{v}$ to train the risk model, focusing on estimating the Value at Risk (VaR) for each instance to aid in risk assessment. This involves identifying high-risk instances through VaR calculations for additional training or correction. Then, it applies the trained risk model to predict labels for the test set $D_{test}$, assigning tentative labels based on the lowest risk as indicated by VaR. Finally, it fine-tunes the classifier using the instances in $D_{test}$ and their predicted labels, with the goal of minimizing misclassification risk as guided by VaR.
@@ 6 Empirical Study
This section evaluates the performance of our proposed approach by an empirical study. It is organized as follows: Subsection. 6.1 describes experimental setup; Subsection. 6.2 presents the evaluation results of risk analysis; Subsection. 6.3 presents the evaluation results of adaptive learning; Subsection. 6.4 presents the evaluation results of sensitivity study; finally, Subsection. 6.5 presents the evaluation results of ablation study.
@@ 6.1 Experimental Setup
We have used two open-sourced benchmark datasets of real H&E stained histopathological breast cancer images, BRACS [6] and BACH. They can be briefly described as follows:
  • 1.
    BRACS. The BReAst Carcinoma Subtyping (BRACS) dataset is a large, annotated collection of 547 Whole-Slide Images (WSIs) and 4539 Regions Of Interest (ROIs) from over 150 patients, aimed at advancing AI for breast cancer characterization. Annotated by three pathologists, the dataset categorizes lesions into three categories and seven subtypes, making it the most comprehensive dataset available for breast cancer subtyping. Table 4 shows the statistics of this dataset.
  • 2.
@@
For misprediction risk analysis, we compare the proposed approach, denoted by MultiRisk with the following alternatives:
  • 1.
@@
For the evaluation of risk-based adaptive learning, we compare our solution with state-of-the-art deep models and approaches for multiclass breast cancer detection, which are described as follows:
  • 1.
    HactNet [33]: HactNet creates Hierarchical Cell-to-Tissue (HACT) graph representations for classifying ROI images and compares results with pathologists’ predictions.
  • 2.
    ScoreNet [40]: ScoreNet is a Transformer-based method that uses scoremix augmentation to achieve top classification performance on BRACS ROIs, comparing favorably with various techniques, including multiple instance learning and GNNs.
  • 3.
    TransPath  [44]: A Transformer-based model for histopathological image classification that utilizes attention mechanisms to effectively capture complex patterns and relationships in image data.
  • 4.
    CTransPath  [45]: The study presents Semantically-Relevant Contrastive Learning (SRCL) to enhance self-supervised learning in medical image analysis by employing a hybrid CTransPath model that combines CNN and Swin Transformer for improved histopathological image tasks.
  • 5.
    SCDA [24]: SCDA (Semantic Concentration for Domain Adaptation) focuses on key features through pair-wise adversarial alignment of prediction distributions, unlike typical distribution alignment methods. This approach is used for domain transfer evaluation.
  • 6.
    TSA [23]: TSA (Transferable Semantic Augmentation) improves classifier adaptation by augmenting source features with random directions from a multivariate normal distribution to reflect semantic changes in the input space, and is applied for domain transfer evaluation.
  • 7.
    CNNs: We also compared our approach with multiple CNNs used to get features for the MultiRisk Approach.
  • 8.
    CCT [27]: The Compact Convolutional Transformer (CCT) merges the strengths of CNNs and transformers, achieving top performance in tasks such as image classification, object detection, and semantic segmentation.
  • 9.
    Ensemble: Ensemble learning for all the CNNs and CCT used to generate risk features for the MultiRisk approach.
  • 10.
@@
As usual, for the evaluation of risk analysis, we use the metric of AUROC; for the evaluation of adaptive learning, we use the metrics of accuracy, precision, recall, F1 and classwise F1.
@@ 6.2 Evaluation Results of Risk Analysis
Figure 5 presents the detailed comparative results for AUROC analysis of MultRisk with SOTA methods. It can be observed that on the original BRACS dataset, the AUROC of the Baseline stands at 73% and the LearnRisk with 74.6%, while ConfidNet, and TrustScore fails to surpass the Baseline’s performance. Remarkably, MultiRisk demonstrates considerable improvement with an AUROC of 78.1%, marking a noteworthy enhancement of 5.1% compared to the Baseline. The results on the BRACS(512 x 512) dataset are similar. The Baseline records an AUROC of 72.9% and LearnRisk with 73.8%, with ConfidNet, and TrustScore falling short of outperforming the Baseline in the multiclass setting. Once again, MultiRisk showcases its effectiveness by achieving an AUROC of 75.6%, with an improvement of 2.7% over the Baseline. These results underscore the efficacy of the MultiRisk approach in enhancing risk analysis performance.
@@
(b) BRACS(512 x 512)
Figure 5: Performance of Risk Analysis on Original and Low Resolution BRACS Respectively
To further evaluate the efficacy of MultRisk, we have also conducted a domain transfer evaluation, where we used the BRACS-BACH four-class dataset. In this experiment, the baseline classifier was trained on BRACS samples but risk model was trained on BACH samples. The detailed evaluation results have been presented in Figure  6. It can be observed that MultiRisk achieved an AUROC of 76.3%, a 2.4% improvement over the Baseline’s 73.9%. While LearnRisk reached 74.3%, both ConfidNet and TrustScore failed to outperform the Baseline. These results highlight MultiRisk’s effectiveness in enhancing risk assessment models in the scenario of domain transfer.
Figure 6: AUROC Evaluation for BRACS-BACH Domain Transfer Task
@@ 6.3 Evaluation Results of Risk-based Adaptive Learning
Table 6: The comparative evaluation results of adaptive learning on original BRACS images
Method Accuracy (%) F1 (%) Precision (%) Recall (%) AUC (%)
Densenet121 60.70 60.65 61.08 60.70 90.84
Resnet50 56.67 55.88 56.20 56.67 87.30
Efficientnet B4 60.53 59.39 58.76 60.53 88.29
CCT 54.39 53.97 55.45 54.39 87.19
Ensemble 59.12 58.36 58.02 59.12 89.63
TransPath 56.27 55.1 55.41 56.27 86.1
CTransPath 57.36 56.75 56.07 57.36 87.14
LearnRisk 62.11 60.05 60.93 61.81 89.30
MultiRisk(Ours) 63.51 61.15 61.80 63.51 91.38
We have presented the detailed comparative results in Table 6 and  7. As shown in Table 6, on the original BRACS dataset, MultiRisk outperforms four deep learning models, including ensemble approach and high-performing architectures like TransPath, CTransPath, and LearnRisk, on all the performance metrics. By enhancing multiclass prediction accuracy through misprediction risk analysis based on the baseline Densenet121, MultiRisk demonstrates robustness and adaptability across datasets.
As shown in Table 7, on the BRACS(512 x 512) dataset, MultiRisk similarly outperforms with 66.34% accuracy, with a 3.01% improvement over the baseline. It also achieves improved performance in terms of F1-score, precision, recall, and AUC, by the margins of 2.57%, 1.72%, 3.01%, and 1.34% respectively. While TransPath and CTransPath had comparable results, MultiRisk consistently secured the highest metrics, highlighting its robustness and efficacy across the different image resolutions.
Table 7: The comparative evaluation results of adaptive learning on BRACS(512 x 512) images
Method Accuracy (%) F1 (%) Precision (%) Recall (%) AUC (%)
Densenet121 63.33 63.41 64.51 63.33 90.73
Resnet50 63.33 62.88 64.56 63.33 89.64
Densenet201 63.16 62.97 63.37 63.16 89.98
CCT 60.70 60.27 60.85 60.70 89.13
Ensemble 62.63 62.70 63.32 62.63 90.74
TransPath 58.17 57.1 57.52 58.17 87.23
CTransPath 60.17 59.7 59.02 60.17 88.7
LearnRisk 65.09 64.23 64.40 64.93 90.10
MultiRisk(Ours) 66.34 65.98 66.23 66.34 92.07
@@
We have also presented the comparative evaluation results of classwise F1-score and weighted F1-score in Table 8. It can be observed that MultiRisk achieves an impressive weighted F1 score of 65.98, surpassing the previous leading transformer-based approach, ScoreNet. Although ScoreNet excels in the flat epithelial atypia class, classical models such as Densenet121, Densenet201, and Resnet50 outperform it in Atypical Ductal Hyperplasia, Usual Ductal Hyperplasia, and Ductal Carcinoma in situ. Notably, MultiRisk improves the F1 score for 6 out of 7 classes in the multiclass setting, demonstrating superior performance in the normal, pathological benign, and Invasive Carcinoma classes compared to the baseline. Specifically, it shows a 6.2% improvement in normal ductal hyperplasia, and 4.41% in flat epithelial atypia. For cancer classes, MultiRisk achieves notable enhancements of 1.6% for ductal carcinoma in situ and 6.73% for invasive carcinoma.
@@
Despite not surpassing SOTA models in the Atypical Ductal Hyperplasia class, this can be attributed to the overall underperformance of non-baseline models in that category. HactNet is widely regarded as the most reliable approach for the BRACS dataset, validated against assessments by real-world pathologists. Notably, MultiRisk outperforms HactNet in nearly every class, achieving an overall improvement of 4.48% in the weighted F1 score.
@@
Table 8: Mean and Standard Deviation evaluation for each class F1 and weighted F1 score using state-of-the-art methods
@@
Densenet121 77.01$\pm$4.5 52.9$\pm$2.6 42.2$\pm$1.9 73.5$\pm$3.5 50.9$\pm$2.5 62.06$\pm$3.1 84.5$\pm$3.5 63.41$\pm$1.5
Densenet201 77.83$\pm$3.9 52.3$\pm$2.3 57.6$\pm$2.7 69.2$\pm$3.3 44.7$\pm$2.0 58.46$\pm$4.2 80.7$\pm$3.6 62.97$\pm$3.1
Resnet50 76.5$\pm$3.7 46.3$\pm$2.5 52.1$\pm$2.6 74.6$\pm$3.5 36.5$\pm$1.6 66.6$\pm$3.2 85.8$\pm$3.7 62.88$\pm$1.4
CCT 73$\pm$4.1 40.6$\pm$2.2 50$\pm$2.4 75.2$\pm$3.5 37.2$\pm$1.7 58.7$\pm$2.8 85.5$\pm$3.6 60.27$\pm$1.7
Ensemble 75.77$\pm$3.3 47.72$\pm$3.0 49.16$\pm$1.8 77.84$\pm$2.8 39.13$\pm$1.8 60.86$\pm$2.7 87.34$\pm$3.1 62.70$\pm$1.9
TransPath 58.8$\pm$2.7 43.4$\pm$2.3 35.2$\pm$4.3 67.2$\pm$6.1 38.6$\pm$1.8 61.7$\pm$1.4 85.4$\pm$1.6 57.1$\pm$2.1
CTransPath 60.0$\pm$1.4 47.1$\pm$2.7 37.9$\pm$2.3 72.9$\pm$3.1 43.3$\pm$3.8 65.9$\pm$1.4 91.0$\pm$2.6 59.7$\pm$1.1
HACT-Net 61.6$\pm$2.1 47.5$\pm$2.9 43.6$\pm$1.9 74.2$\pm$1.4 40.4$\pm$2.5 66.4$\pm$2.6 88.4$\pm$0.2 61.5$\pm$0.9
ScoreNet 64.3$\pm$1.5 54.0$\pm$2.2 45.3$\pm$3.4 78.1$\pm$2.8 46.7$\pm$1.0 62.9$\pm$2.0 91.0$\pm$1.4 64.4$\pm$0.9
LearnRisk 77.71$\pm$4.4 53.98$\pm$2.6 46.97$\pm$2.1 76.92$\pm$3.1 44.11$\pm$2.6 61.45$\pm$1.8 88.46$\pm$1.3 64.23$\pm$0.9
MultiRisk(Ours) 78.30$\pm$4.3 54.71$\pm$2.6 48.40$\pm$2.3 77.91$\pm$3.9 47.62$\pm$2.1 63.66$\pm$2.1 91.23$\pm$2.7 65.98$\pm$1.0
To further validate the efficacy of MultiRisk, we have also conducted an empirical study on the domain-transfer BRACS-BACH dataset where training images come from BRACS but validation and test images come from BACH. Besides the existing leading deep models, we also compare MultiRisk with an ensemble of these models, LearnRisk and two SOTA domain transfer learning approaches, SCDA and TSA.
Table 9: Accuracy, F1-Score, Precision, Recall, and AUC Evaluation for BRACS-BACH Domain Transfer Task
Model Accuracy (%) Precision (%) Recall (%) AUC (%)
Efficientnet_B4 77.50 79.08 77.50 92.35
Densenet121 73.75 74.17 73.75 88.58
CCT 70.00 71.35 70.00 89.38
Resnet50 67.50 67.93 67.50 87.65
Ensemble 70.00 72.65 70.00 89.54
SCDA 78.91 79.45 78.91 89.74
TSA 72.64 72.94 72.64 86.76
LearnRisk 77.50 79.25 77.50 92.93
MultiRisk(Ours) 80.00 81.90 80.00 93.23
Table 10: F1 Score Evaluation for BRACS BACH Domain Transfer Task
Model Normal Benign InSitu Invasive Weighted F1
EfficientnetB4 86.48±4.2 60.46±3.5 78.04±4.3 87.17±4.5 78.04±1.0
Resnet50 72.34±4.5 29.41±2.4 69.76±4.0 77.77±3.5 67.33±1.9
Densenet121 78.04±4.2 55.00±3.0 80.95±4.4 81.08±3.4 73.77±1.0
CCT 87.17±4.5 47.36±3.1 65.21±4.4 81.08±3.5 70.21±1.5
Ensemble 86.48±±4.2 51.16±3.2 68.18±4.0 83.33±3.5 70.84±1.5
SCDA 86.51±m4.2 60.63±3.3 80.50±4.1 87.12±m3.9 78.69±1.0
TSA 86.45±3.1 65.92±3.9 71.05±2.6 70.64±4.7 72.01±1.3
LearnRisk 83.33±4.1 60.46±3.5 80.95±4.3 87.17±4.5 77.98±1.0
MultiRisk(Ours) 88.88±4.1 65.11±3.5 80.95±4.3 87.17±4.5 80.53±0.9
We have presented the overall comparative evaluation results in Table 9. It can be observed that MultiRisk outperforms its alternatives on all the metrics, underscoring its efficacy in domain-transfer scenarios. We have also presented the detailed class-wise F1 scores in Table 10. It can be observed that MultiRisk outperforms the baseline (EfficientNet-B4), with improvements of 2.4% for Normal, 4.64% for Benign, and 2.9% for InSitu, while showing no considerable change for Invasive. Although SCDA also outperforms the deep models and even LearnRisk, TSA shows limited efficacy compared with them, because ideally TSA was built for large amount of data from all the domains, and in this scenario due to the limited number of validation and test images compared to the cross-domain training images its efficacy is limited.

### v2

@@ 4.2 Attention-based Risk Model Construction
The existing binary model supposes risk features are mutually independent. However, in real scenarios, risk features are to a large extent correlated, which is especially true in the multi-class scenario. Therefore, we present a new dynamic risk model with attention-based weighting for multi-class risk analysis. The attention mechanism dynamically adjusts the importance of risk features based on the context of each class, improving multiclass prediction by weighting features according to their relevance [77, 78].
@@
In a multi-class prediction task, consider a sample $d_{i}$ that can belong to one of four classes: A, B, C, or D. In the learnable Gaussian model, each class has a set of mean ($\mu_{A},\mu_{B},\mu_{C},\mu_{D}$) and variance ($\sigma_{A}^{2},\sigma_{B}^{2},\sigma_{C}^{2},\sigma_{D}^{2}$) parameters that are learned from the risk features. These parameters are used to compute the Gaussian weights, however, even with these learnable parameters, the model still uses fixed, class-specific parameters that do not change dynamically based on the sample content. In contrast, the attention mechanism assigns dynamic weights based on the risk features. For example, for image $d_{i}$, the attention mechanism might assign $\omega_{A}=0.6$, $\omega_{B}=0.1$, $\omega_{C}=0.2$, and $\omega_{D}=0.1$, emphasizing class A more than the others based on the sample’s specific features. This dynamic adjustment allows the attention mechanism to focus on the most relevant classes for each input sample, whereas the Gaussian model is constrained by pre-learned, static parameters.
@@ 4.3 Risk Model Training
Class imbalance is common in most histopathology image datasets, as shown in Tables 5 and 6, particularly in the training data. To address this issue, we propose to neutralize class bias in expectation and variance for balanced learning and prediction. The neutralization operation is formally defined by Equation 15, and 16 as follows:
@@
7:Maximum number of training iterations NN
@@
where $\gamma$ denotes the risk score of the particular classes of corresponding samples. Considering this scenario, the posterior probability $p_{ij,c}$ of the image $(d_{i},c_{i})$ is ranked higher than $(d_{j},c_{j})$ is given by:
@@
In a breast cancer subtype prediction task with four classes (benign, malignant, in situ, and invasive), consider two images, $d_{i}$ and $d_{j}$, with predicted class probability vectors $d_{i}=[0.2,0.6,0.1,0.1]$ and $d_{j}=[0.3,0.4,0.2,0.1]$. Performing pairwise comparisons, $d_{i}$ wins against $d_{j}$ in the malignant class (1 win) but loses in the benign and in situ classes (0 wins each), while both images tie in the invasive class. The total win count results in $W_{i}=1$ for $d_{i}$ and $W_{j}=2$ for $d_{j}$. Consequently, $d_{j}$ is ranked higher due to its superior win count, demonstrating the effectiveness of the voting-based mechanism for class-based image ranking.
@@
We also introduced dropout during training to randomly set a percentage of input units to zero. This prevents co-adaptation and promotes the learning of more robust features for each class. Dropout is inactive during evaluation, with activations scaled for consistency. In multiclass prediction, dropout helps prevent overfitting by ensuring that the model does not overly rely on specific features. We set the dropout rate to 0.5 and fine-tuned it for optimal performance.
@@ 5 Solution for MultiRisk Based Adaptive Training
For the task of breast cancer subtype classifcation, neural networks using softmax often produce overly confident predictions, which can mislead training and impair generalization [73]. Class imbalance can further degrade performance on minority classes [74], and risk estimation may fail to accurately reflect misprediction likelihood, particularly in complex multiclass scenarios [35].
@@
  • Traditional Pre-Training Phase: During this initial phase, a deep model undergoes training using conventional methods, relying on labeled training data for breast cancer subtypes. In our method, we use four CNN deep models and train all of them on the given workload with arguably the best parameters. The aim is to try to achieve the best test efficacy for each model.
  • Risk Based Adaptive Training Phase: This phase aims to enhance model performance by focusing on high-risk instances and managing prediction uncertainties. This framework integrates temperature scaling to reduce overconfidence and Value at Risk (VaR) to assess misprediction risks. The goal is to improve the robustness and accuracy of the prediction model. The step-wise process is given below:
Algorithm 3 Risk-Based Adaptive Training for Breast Cancer Subtype Prediction
Input: A workload of breast cancer subtype histopathological images $D$ consisting of $D_{s}$, $D_{v}$, $D_{t}$ and initial classifier $g(\omega)$;
@@
[…]
2:for $k=0$ to m1m-1 do
@@
In the second phase, the algorithm first uses the validation set $D_{v}$ to train the risk model, focusing on estimating the Value at Risk (VaR) for each instance to aid in risk assessment. This involves identifying high-risk instances through VaR calculations for additional training or correction. Then, it applies the trained risk model to predict labels for the test set $D_{test}$, assigning tentative labels based on the lowest risk as indicated by VaR. Finally, it fine-tunes the classifier using the instances in $D_{test}$ and their predicted labels, with the goal of minimizing misprediction risk as guided by VaR.
@@ 6 Empirical Study
This section evaluates the performance of our proposed approach by an empirical study. It is organized as follows: Subsection. 6.1 describes experimental setup; Subsection. 6.2 presents the evaluation results of risk analysis; Subsection. 6.3 presents the evaluation results of adaptive learning; Subsection. 6.4 presents the evaluation results of sensitivity analysis; Subsection. 6.5 presents the evaluation results of ablation study; Subsection. 6.6 presents the evaluation results of generalization on other type of cancers; and finally, Subsection. 6.7 presents the presents the discussion on computational efficiency and scalability of our method;
@@ 6.1 Experimental Setup
We have used two open-sourced benchmark datasets of real H&E stained histopathological breast cancer images, BRACS [10] and BACH [9]. They can be briefly described as follows:
  • BRACS. The BReAst Carcinoma Sub-typing (BRACS) dataset is a large, annotated collection of 547 Whole-Slide Images (WSIs) and 4539 Regions Of Interest (ROIs) from over 150 patients, aimed at advancing AI for breast cancer characterization. Annotated by three pathologists, the dataset categorizes lesions into three categories and seven subtypes, making it the most comprehensive dataset available for breast cancer sub-typing. Table 5 shows the statistics of this dataset.
@@
Furthermore to demonstrate the generalizability of our approach beyond breast cancer, we also conducted experiments on two additional cancer histopathology datasets: LC25000 [98], which contains 25,000 images of lung and colon cancer across five classes including lung adenocarcinoma, lung squamous cell carcinoma, colon adenocarcinoma, lung benign, and colon benign tissues; and LungHist700 [99], comprising 700 images with three classes focusing on lung adenocarcinoma, lung squamous cell carcinoma, and normal lung tissue.
@@
For misprediction risk analysis, we compare the proposed approach, denoted by MultiRisk with the following SOTA approaches:
@@
For the evaluation of risk-based adaptive learning, we compare our solution with state-of-the-art methods from the perspective of multiple categories, inlcuding backbones, SOTA multiclass breast cancer and histopathological image classification, domain adaptation, and uncertainty estimation-based training:
  • Backbones:
    • CNNs: Multiple CNN architectures were used to extract features for the MultiRisk approach.
    • CCT [63]: The Compact Convolutional Transformer merges the strengths of CNNs and transformers, achieving top performance in tasks such as image classification, object detection, and semantic segmentation.
  • SOTA Breast Cancer & Histopathological Image Prediction:
    • HactNet [61]: Creates Hierarchical Cell-to-Tissue (HACT) graph representations for predicting subtypes from ROI images and compares results with pathologists’ predictions.
    • ScoreNet [32]: Transformer-based method using scoremix augmentation to achieve top prediction performance on BRACS ROIs, outperforming techniques like multiple instance learning and GNNs.
    • TransPath [91]: Transformer-based model utilizing attention mechanisms to effectively capture complex patterns and relationships in histopathological image data.
    • CTransPath [92]: Employs Semantically-Relevant Contrastive Learning (SRCL) in a hybrid CNN–Swin Transformer model to improve histopathological image tasks.
    • TransMIL [96]: A transformer-based Multiple Instance Learning framework tailored for whole-slide image classification in pathology.
    • CLAM [97]: Clustering-constrained Attention Multiple Instance Learning framework that provides interpretable histopathology classification.
  • Domain Adaptation:
    • SCDA [44]: Semantic Concentration for Domain Adaptation focuses on key features through pair-wise adversarial alignment of prediction distributions.
    • TSA [45]: Transferable Semantic Augmentation improves classifier adaptation by augmenting source features with semantically-relevant random directions.
    • DANN [95]: Domain-Adversarial Neural Network using gradient reversal to learn domain-invariant representations for improved generalization.
  • Uncertainty Estimation(Risk)–Based Training:
    • MCdropout [93]: Monte Carlo dropout is applied during inference to estimate predictive uncertainty by sampling multiple stochastic forward passes.
    • Deep Ensemble [94]: Combines predictions from multiple independently trained models to capture both model and data uncertainty for more robust predictions. In our case three CNNs combined with CCT
@@
  • MultiRisk(Ours): Uses misprediction risk analysis designed for subtype prediction and adaptive training to enhance the performance of state-of-the-art approaches.
Furthermore, we conducted experiments using large language model–based vision-language models, using the concept of CLIP [104], and DeepSeek-VL [100] in a supervised manner, on the datasets mentioned above. The performance of these vision–language models was very low on this task, and the gap compared to other state-of-the-art methods was substantial. For this reason, we do not include their results in the main experiments. However, to demonstrate the generalizability of our proposed approach, we implemented our adaptive learning method on these models. With this integration, we were able to achieve significantly better performance. These results are presented in Section 6.6. For the evaluation of risk analysis, we use the AUROC metric; for adaptive learning, we report the F1 score both per class and as a weighted average, along with AUC.
@@ 6.2 Evaluation Results of Risk Analysis
Figure 5 presents the detailed comparative results for AUROC analysis of MultRisk with SOTA methods On Original BRACS dataset, BRACS(512 x 512) variant, and the BRACS-BACH domain transfer dataset. It can be observed that on the original BRACS dataset in figure 5 (a), the AUROC of the Baseline stands at 73% and the LearnRisk with 74.6%, while ConfidNet, and TrustScore fails to surpass the Baseline’s performance. Remarkably, MultiRisk demonstrates considerable improvement with an AUROC of 78.1%, marking a noteworthy enhancement of 5.1% compared to the Baseline. The results on the BRACS(512 x 512) dataset in figure 5 (b) are similar. The Baseline records an AUROC of 72.9% and LearnRisk with 73.8%, with ConfidNet, and TrustScore falling short of outperforming the Baseline in the multiclass setting. Once again, MultiRisk showcases its effectiveness by achieving an AUROC of 75.6%, with an improvement of 2.7% over the Baseline. Furthermore, for BRACS-BACH domain transfer four-class dataset, the baseline classifier was trained on BRACS samples but risk model was trained on BACH samples. The detailed evaluation in Figure 5 (c), It can be observed that MultiRisk achieved an AUROC of 76.3%, a 2.4% improvement over the Baseline’s 73.9%. While LearnRisk reached 74.3%, both ConfidNet and TrustScore failed to outperform the Baseline. These results highlight MultiRisk’s effectiveness in enhancing risk assessment models in the scenario of domain transfer.
These results underscore the efficacy of the MultiRisk approach in enhancing risk analysis performance.
@@
(b) BRACS (512 × 512)
(c) BRACS-BACH
Figure 5: Performance of Risk Analysis on (a) Original BRACS, (b) BRACS(512×512), and (c) BRACS-BACH datasets.
@@ 6.3 Evaluation Results of Risk-based Adaptive Learning
We present detailed comparative results in Table 7. On both the original BRACS and BRACS (512×512) datasets, results are reported for backbone methods, state-of-the-art (SOTA) breast cancer subtype and histopathological image prediction methods, as well as uncertainty estimation-based Learning techniques. For the domain transfer task, three SOTA domain adaptation methods are included. Across all three datasets and four categories of SOTA models, MultiRisk consistently outperforms the alternatives. By enhancing subtype prediction accuracy through misprediction risk analysis applied on baselines such as DenseNet121 and EfficientNetB4, MultiRisk demonstrates robustness and adaptability across datasets.
As shown in Table 7, MultiRisk achieves an F1-score of 61.15% and an AUC of 91.38% on the original BRACS dataset. Similarly, on the BRACS (512×512) dataset, it attains an F1-score of 65.98% and an AUC of 92.07%. While some methods report comparable results, MultiRisk consistently secures the highest metrics, highlighting its robustness and efficacy across different image resolutions. Furthermore, for the domain transfer task (BRACS-BACH), MultiRisk outperforms alternative approaches such as TSA, SCDA, and DANN in terms of both F1-score and AUC, underscoring its effectiveness in domain-transfer scenarios.
Table 7: Performance comparison (F1-score % & AUC %) on BRACS (original), BRACS(512x512), and BRACS→BACH (domain shift). Best results are presented in bold. (mean and std across 5 runs)
[…]
F1 AUC F1 AUC F1 AUC
[…]
@@
Furthermore, BRACS is considered the benchmark data for breast cancer subtype prediction for which classwise F1 scores with weighted F1 are presented in state-of-the-art breast cancer subtype prediction methods like HACT-net, and ScoreNet. We have also presented the comparative evaluation results of classwise F1-score and weighted F1-score in Table 8 for the backbone methods, state-of-the-art (SOTA) breast cancer subtype and histopathological image prediction methods, as well as uncertainty estimation-based learning techniques. It can be observed that MultiRisk achieves an impressive weighted F1 score of 65.98, surpassing the previous leading transformer-based approach, ScoreNet. Although ScoreNet excels in the flat epithelial atypia class, classical models such as Densenet-121, Densenet-201, and Resnet-50 outperform it in Atypical Ductal Hyperplasia, Usual Ductal Hyperplasia, and the histpathological image prediction method TransMIL outperforms it in Ductal Carcinoma in situ. Notably, MultiRisk improves the F1 score for 6 out of 7 classes considering the baseline in the subtype setting, demonstrating superior performance in the normal, pathological benign, and Invasive Carcinoma classes compared to the baseline. Specifically, it shows a 6.2% improvement in normal ductal hyperplasia, and 4.41% in flat epithelial atypia. For cancer classes, MultiRisk achieves notable enhancements of 1.6% for ductal carcinoma in situ and 6.73% for invasive carcinoma.
@@
Despite not surpassing SOTA models in the Atypical Ductal Hyperplasia subtype, this can be attributed to the overall under-performance of non-baseline models in that category. HactNet is widely regarded as the most reliable approach for the BRACS dataset, validated against assessments by real-world pathologists. Notably, MultiRisk outperforms HactNet in nearly every class, achieving an overall improvement of 4.48% in the weighted F1 score.
@@
Table 8: Mean and Standard Deviation evaluation for each class F1 and weighted F1 score using state-of-the-art methods for Benchmark BRACS Dataset(mean and std across 5 runs)
@@
[…]